OpenStudy (anonymous):

In the supermarket, there is a scale that measures the weight of fruit and vegetables. The needle is 24 cm long and it rotated 96° when Charlie placed four tomatoes on it. How far did the needle travel as it rotated? @M4thM1nd

OpenStudy (anonymous):

@M4thM1nd Please just leave answer I got to go dinner time and relaxation got to go

OpenStudy (anonymous):

When we work with angles in radians we have this relationship: l = r*theta, where l is the arc length, r is the radius and theta is the angle in radians From this question we have r = 24 cm = 0.24 m theta = 96º = (96º * 2*pi)/(360º) = 8*pi/15 So... the traveled by the end point of the needle is: l = r*theta l = (0.24 m)*(8*pi/15) l = 0.402124 m l = 40.2124 cm

OpenStudy (anonymous):

Or l = 64*pi/5 cm